An electron experiences a force of 8.7 x 10 -12 N. What is the intensity of the electric field?
1
Expert's answer
2020-02-24T11:02:51-0500


A force acts on an electron in an electric field

"F=e \cdot E"

Where will we write

"E=\frac{F}{e}=\frac{8.7 \cdot 10^{-12}}{1.6 \cdot 10^{-19}}=5.438\cdot10^{7}V/m"